### v1

# OSS License Identification at Scale: A Comprehensive Dataset Using World of Code

Abstract.
The proliferation of open source software (OSS) has led to a complex landscape of licensing practices, making accurate license identification crucial for legal and compliance purposes. This study presents a comprehensive analysis of OSS licenses using the World of Code (WoC) infrastructure. We employ an exhaustive approach, scanning all files containing “license” in their filepath, and apply the winnowing algorithm for robust text matching. Our method identifies and matches over 5.5 million distinct license blobs across millions of OSS projects, creating a detailed project-to-license (P2L) map. We verify the accuracy of our approach through stratified sampling and manual review, achieving a final accuracy of 92.08%, with precision of 87.14%, recall of 95.45%, and an F1 score of 91.11%. This work enhances the understanding of OSS licensing practices and provides a valuable resource for developers, researchers, and legal professionals. Future work will expand the scope of license detection to include code files and references to licenses in project documentation.
Software License, Open Source Software, Open Source License, World of Code
ccs: Software and its engineering Software creation and managementccs: General and reference Empirical studies
1. Introduction
In recent years, the open source software (OSS) landscape has expanded rapidly, leading to a diverse ecosystem of projects with varying licensing practices. The identification and analysis of these licenses are crucial for understanding the legal frameworks that govern OSS distribution and usage. Previous studies have explored specific subsets of OSS licensing; however, a comprehensive analysis across the entire open source landscape has been lacking (Reid and Mockus, 2023). In this study, we present a comprehensive dataset of open source projects and their licenses, utilizing the World of Code (WoC) infrastructure to perform an exhaustive scan of the OSS landscape. Specifically, we target any file with “license” in its filepath, ensuring a broad coverage of licensing practices.
To enhance the accuracy of license detection, we employ the winnowing algorithm, a robust technique known for its effectiveness in text pattern matching, even in cases where the text is embedded or has undergone slight modifications (Serafini and Zacchiroli, 2022). This approach allows us to accurately identify and match millions of distinct license blobs with known licenses, thereby creating a comprehensive database that captures the full diversity of open source licensing. By doing so, our research broadens the scope of existing literature, providing a cross-platform understanding of licensing practices across the OSS ecosystem.
This dataset and the insights derived from it will serve as a valuable resource for developers, researchers, and legal professionals engaged in the open source community, enabling a deeper understanding of the legal structures underpinning OSS.
2. Related Work and Contributions
2.0.1. Comprehensive Identification and Matching of License Blobs
Previous studies like Wu et al. (Wu et al., 2024) and Xu et al. (Xu et al., 2023) focus on explicit license declarations in metadata files, while others, such as Feng et al. (Feng et al., 2019), use static analysis to detect embedded license texts in binaries. However, these methods may miss licenses not explicitly declared or located in unconventional directories. Unlike these approaches, our research adopts a more exhaustive strategy by scanning the entire OSS landscape for any file with “license” in its filepath. This ensures that both standard and non-standard files containing licensing information are included.
Moreover, we employ the winnowing algorithm, a robust method for matching license texts to known licenses, enhancing the accuracy of detecting both partial and full matches, even when the text is embedded or slightly modified.
2.0.2. Scale and Scope of Analysis
Previous research often focuses on specific platforms (e.g., GitHub), package managers (e.g., NPM), or specific license types (e.g., OSI-approved licenses), limiting their scale and comprehensiveness. Our study broadens this scope by analyzing the entire open source landscape, providing a cross-platform understanding of licensing practices. By identifying millions of distinct license blobs and matching them with known licenses, we create a comprehensive database that captures the full diversity of open source licensing.
3. Methodology
3.1. World of Code Infrastructure
World of Code (WoC) [^1] https://worldofcode.org is an infrastructure designed to cross-reference source code changes across the entire FLOSS community, enabling sampling, measurement, and analysis across software ecosystems (Ma et al., 2019, 2021). It functions as a software analysis pipeline, handling data discovery, retrieval, storage, updates, and transformations for downstream tasks (Ma et al., 2021).
WoC offers maps connecting git objects and metadata (e.g., commits, blobs, authors) and higher-level maps like project-to-author connections, author aliasing (Fry et al., 2020), and project deforking (Mockus et al., 2020). We use WoC to identify all license blobs and their associated projects [^2] Version V, latest at the time of this study., employing the concept of deforked projects (Mockus et al., 2020) to avoid biases from forks and duplicates. “Project” refers to these deforked projects unless noted otherwise.
3.2. Finding License Blobs
Our goal is to identify the license(s) applicable to each open source project. We start by using the project-to-file maps (P2f) in WoC to list all files in each project, specifically searching for those with “license” in their filepath. This approach resulted in over 10 million distinct potential license blobs.
Since there are relatively few known licenses, we anticipate that most of these blobs are similar licenses with minor variations, such as differences in whitespace, formatting, or non-essential information. The main challenge is matching these varied license blobs to known licenses.
We use licenses from the Open Source Initiative [^3] https://github.com/OpenSourceOrg/licenses and the Software Package Data Exchange (SPDX) [^4] https://github.com/spdx/license-list-data, which include 103 and 635 licenses, respectively. To match the 10 million potential license blobs with these known licenses, we apply winnowing, an efficient local fingerprinting algorithm (Schleimer et al., 2003).
Winnowing is a document fingerprinting technique often used in plagiarism detection. It generates fingerprints by sliding a window over hashed words in a document, selecting the smallest hash value in each window. This reduces the data needed for document representation, enabling faster and more memory-efficient comparisons while maintaining accuracy.
Using winnowing, we matched over 7 million potential license blobs to known licenses (see Table 1). We assess the reliability of these matches by calculating a matching score, defined as the number of shared winnowing signatures divided by the total winnowing signatures between two files. This score, as shown in Equation 1, measures the similarity between the potential license blob and the known license, helping to verify the match’s accuracy.
(1) $S=\frac{c(A\cap B)}{c(A\cup B)}$
$S$::
The matching score.
$A$::
The set of signatures in document A.
$B$::
The set of signatures in document B.
$c(X)$::
The count function that returns the number of elements in set $X$.
Table 1. Potential License Blobs Matching Scores
Count Percentage (Relative to) Percentage (Overall)
Potential Blobs 10,093,268 100% 100%
Winnowing 9,794,559 97% (Potential Blobs) 97%
Matched 7,167,046 73.2% (Winnowing) 71%
S ¡= 0.2 795,532 11.1% (Matched) 7.9%
0.2 ¡ S ¡= 0.4 239,091 3.3% (Matched) 2.4%
0.4 ¡ S ¡= 0.6 264,667 3.7% (Matched) 2.6%
0.6 ¡ S ¡= 0.8 435,283 6.1% (Matched) 4.3%
0.8 ¡ S ¡= 1 5,432,473 75.8% (Matched) 53.8%
We categorized matching scores into five groups: below 20%, 20-40%, 40-60%, 60-80%, and above 80%. As shown in Table 1, 97% of blobs generated winnowing signatures. We randomly sampled 30 blobs from the 3% that did not and manually confirmed they had no meaningful content. Of the 9.7 million blobs, 73% matched a known license (sharing at least one winnowing signature), with 75% of these matches scoring above 80%.
To assess match reliability, we sampled 20 blobs from each score group and manually compared them to the known license using ‘diff‘. Given the manual nature of the verification process, choosing 20 samples for each bucket provides a manageable workload while still offering a sufficient range of data to detect patterns and inconsistencies. Our investigation revealed that matches in buckets with scores below 80% were not reliable enough, showing meaningful differences.
We then focused on scores above 80% and conducted another stratified sampling based on score range (80-85, 85-90, 90-95, 95-100) and the number of winnowing signatures (above/below 100). In these eight groups, 20 matches were sampled and reviewed. Our investigation showed that almost all cases were reliable. The differences fell into three main categories: 1. Identical content with different formatting, 2. Identical content with non-license text, and 3. Identical content with additional clauses.
The second category was acceptable, as we don’t claim a blob contains only the matched license. However, the third, with additional clauses, was concerning as it could alter the license’s nature. Detailed results are in Table 2.
Table 2. Matching Score Samples
Signatures Score Total Count (%) Sample Size Group 1 Group 2 Group 3
¡=100 80-85 85,294 (1.6%) 20 17 3 0
85-90 150,046 (2.8%) 20 17 3 0
90-95 197,875 (3.6%) 20 20 0 0
95-100 4,502,264 (82.9%) 20 20 0 0
¿100 80-85 67,235 (1.2%) 20 10 9 1
85-90 52,894 (1%) 20 17 2 1
90-95 60,583 (1.1%) 20 18 2 0
95-100 316,282 (5.8%) 20 20 0 0
We observed only two mismatches: one in the 80-85% range and one in the 85-90% range (both in the over 100 signatures group).
Based on this, we determined that setting the threshold at 85% ensures reliable license identification. Above this threshold, critical mismatches—where additional clauses could alter the license—are extremely rare. Since over 90% of identified blobs had fewer than 100 winnowing signatures, the 85% threshold balances comprehensiveness and precision, capturing most valid matches while minimizing misleading results. This approach aligns with prior research emphasizing high similarity thresholds to reduce false positives in textual matching (e.g., (Kapitsaki et al., 2017)). As a result, 5,294,666 distinct blobs were matched with a known license.
For the remaining 2.5 million potential blobs with no matches, we randomly sampled 30 and manually investigated them. Only 5 contained license-related content, either mentioning a license name or linking to a license URL. The other 25 were unrelated to licenses.
3.3. Project to License Map
To create the project-to-license (P2L) map, we use the 5.5 million matched license blobs and join them with WoC’s blob-to-time project (b2tP) map, which links blobs to the projects they were committed to, along with commit timestamps. This produces a table mapping each project to a known license and the time of the commit.
However, a blob’s presence in a project’s latest status can’t be confirmed solely from commit history, as it might have been removed later. To address this, we use the project-to-last-commit (P2lc) and tree-to-objects (t2all) maps from WoC. The P2lc map links projects to their last commit at the time of the latest WoC update (Version V), allowing us to retrieve the list of all blobs in a project’s current state by joining P2lc, c2dat (commit-to-tree), and t2all maps. This method not only provides all the times at which a blob was committed to a project but also verifies whether it still exists in the project.
The final table is saved as a semicolon-separated file containing three fields555 For more information on accessing this data, please visit https://github.com/woc-hack/tutorial:
$Project\_ID;License;Commit\_Timestamp$
3.4. P2L Verification
For the Project-to-License (P2L) verification, we initially sampled 1,000 projects from approximately 130 million to evaluate the effectiveness of our license assignment methodology. This sample size was chosen to provide a statistically significant subset for manual verification while balancing the need for reliability with the practical constraints of manual inspection.
We stratified the sample into three groups: 1. Projects with matched licenses, where our automated process successfully matched license blobs to known licenses, 2. Projects with license blobs but no matched licenses, where license blobs were identified, but no matching known license could be confirmed, and 3. Projects without any license blobs, where no license blobs were detected during the automated search.
This sampling approach was designed to cover a wide range of license detection scenarios, ensuring a comprehensive evaluation. Graduate students manually reviewed the sampled projects as part of a class assignment, focusing on verifying the license information. Out of the 1,000 sampled projects, we received meaningful responses for 580 projects, distributed as follows: 291 with matched licenses, 139 with license blobs but no matches, and 150 without any license blobs. The results are presented in Table 3(a).
Table 3. License Detection Confusion Matrix
(a) Initial
License No Lic.
Matched 210 81
Not Matched 22 267
Accuracy 82.24%
Precision 72.16%
Recall 90.52%
F1 Score 80.31%
(b) Adjusted
License No Lic.
210 31
22 267
90.00%
87.14%
90.52%
88.79%
(c) Refined
License No Lic.
210 31
10 267
92.08%
87.14%
95.45%
91.11%
Our license detection method demonstrated reasonable performance with an initial accuracy of 82.24%, precision of 72.16%, recall of 90.52%, and an F1 score of 80.31%.
However, several factors must be considered when interpreting these results: 1. Of the 81 projects identified as having matched licenses, 39 no longer exist on GitHub, preventing license verification, and 2. In 11 projects, the license was absent in the latest status, which does not necessarily indicate a false positive, as the license could have been removed after an earlier commit.
After excluding these cases, we are left with 31 false positives. Adjusting for these, our revised performance metrics show significant improvement: accuracy increases to 90.00%, precision to 87.14%, recall remains at 90.52%, and the F1 score rises to 88.79%.
For the 22 false negatives (where licenses were not detected), further investigation revealed that only 10 had a missed license blob, which was matched but fell slightly below our 85% threshold. The remaining 12 projects only referenced a license (e.g., in the README) without including the actual license file in the repository, so they were not expected to be matched by our method.
By excluding these 12 false negatives, which fall outside our method’s intended scope, we can more accurately assess its performance. The recalculated metrics show an accuracy of 92.08%, precision of 87.14%, recall of 95.45%, and an F1 score of 91.11% (see Table 3(c)).
4. Limitations and Future Work
4.0.1. License Files
Currently, we search for license information only in files containing “license” in their path, such as LICENSE or LICENSE.txt, or files within a “license” sub-directory. Future work could expand this search to include other files, particularly code files with license information commented at the top. Parsing all code blobs to identify license information within them would be a valuable enhancement.
4.0.2. License Pointers
Our current method focuses on file contents to identify licenses, but some projects only reference a license by name or URL without copying the full text. These references aren’t included in our license list. A future improvement would be developing an algorithm to reliably parse README files, project documentation, and other potential sources to identify such license pointers.